# Exact Chiral Symmetries of 3+1D Hamiltonian Lattice Fermions

### v2

@@ I Introduction
Regulating chiral gauge theories like the standard model on the lattice has been a long standing problem. As well as offering a route for extracting more precise predictions from these theories, this has been highlighted as a deep theoretical problem thanks to early no-go results like the Nielsen-Ninomiya theorem [1, 2]. These prove that lattice theories with certain features always come with fermions of both handedness, a phenomenon known as fermion doubling. This problem already exists for fermions with chiral global symmetries, and can be understood as a consequence of ’t Hooft anomaly matching [3]. In particular, a system with an on-site global symmetry, meaning one which does not mix degrees of freedom at separate spacetime points, must be free of ’t Hooft anomalies in the infrared (IR). This rules out, for example, a 3+1D lattice model with an on-site $U(1)$ global symmetry giving rise to a single charged Weyl fermion at low energies.
There have been several approaches so far to circumvent the fermion doubling problem. A simple strategy is to try to “gap out” the unwanted fermions by introducing carefully chosen mass terms. Examples of this construction are Wilson fermions [4] and Kogut-Susskind fermions [5]. These fermions are commonly used in lattice simulations of QCD, whose gauge symmetry is vector-like (non-chiral) and can be realized in an on-site fashion, allowing coupling to gauge fields.
Because of anomaly matching, anomalous chiral flavor symmetries of these models are explicitly broken by the mass terms. This puts the fermions in danger of being fine tuned, unless there is enough remnant symmetry to protect them from obtaining other mass terms (i.e. to prevent additive mass renormalization). This usually comes in the form of discrete crystalline symmetries of the lattice model such as translation, which can act chirally. This mechanism also stabilizes gapless fermions in condensed matter systems such as Weyl semimetals [6, 7, 8, 9, 10].
Another method to stabilize such fermions is to use more general not-on-site symmetries, which can give rise to continuous chiral symmetries. An early example is the Ginsparg-Wilson symmetry of a Euclidean lattice fermion [11, 12] which infinitesimally takes takes the form
@@
where $a$ is the lattice spacing, and $D$ is a finite difference operator which is a discretization of the spacetime Dirac operator. The operator $(D\psi)(x)$ is a (linear) function of $\psi$ at nearby spacetime lattice points, which makes it not-on-site, whereas an on-site symmetry would be a function of $\psi(x)$ alone (and therefore anomaly-free). When $a\to 0$ this becomes the continuum axial symmetry of a Dirac fermion. It thus protects the fermion from gaining a mass. To emphasize that this axial symmetry is not just emergent, but that it is forced by the UV Ginsparg-Wilson symmetry, we call it an “emanant symmetry” after [13].
In this paper, we will construct Hamiltonian models with minimal number of fermions at low energy, transforming anomalously under not-on-site symmetries. In particular, we construct three main models
@@
    a 3+1D model with a single Weyl fermion having an emanant $U(1)$ chiral symmetry
@@
    a 3+1D model with a pair of Weyl fermions having an emanant $SU(2)$ chiral symmetry with Witten’s global $SU(2)$ anomaly [14]
  3. 3.
    a 2+1D model with a single Dirac fermion and an emanant $U(1)\rtimes T$ symmetry with the parity anomaly
We emphasize that these models have the minimal number of fermions for their emanant anomalous symmetries, and that each is protected from gaining a mass—they are not fine tuned. Other Hamiltonian families of models of fermions enjoying Ginsparg-Wilson symmetries exist, such as overlap fermions [15, 16, 17, 18]. As far as we know, our constructions give the first with these minimal representations and emanant continuous symmetry groups.
From the condensed matter point of view, these are models of Weyl semimetals [6, 7]. Using band theory, and especially the Bogoliubov-de-Gennes (BdG) formalism, we are able to study a large class of transformations generalizing the Ginsparg-Wilson symmetry (1) which are linear in the fermion operators, which we leverage to produce the models. We can also derive some no-go results in this formalism, constraining what these not-on-site symmetries must look like for certain anomalous symmetries to emanate from them, making contact with other no-go theorems generalizing Nielsen-Ninomiya such as in [19].
@@ II Symmetry-protected single Weyl fermion in 3+1d
Here we build a (time-reversal broken) 3+1D tight-binding model with finite-range hopping and a single Weyl node at crystalline momentum $\mathbf{k}=\mathbf{0}$ that is protected by a finite-range non-on-site chiral symmetry. We start with a two-band model (two fermion species per site, which we think of as spin $s\in\{\uparrow,\downarrow\}$) on a cubic lattice, known as a magnetic Weyl semimetal, described by the second-quantized Hamiltonian
@@
with $m(\mathbf{k})=2-\cos k_{x}-\cos k_{y}$, and $\boldsymbol{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are Pauli matrices acting on the spin degree of freedom. It hosts two Weyl fermions at low energy, seen by linearizing the Hamiltonian are the node at momentum $\mathbf{k}_{1}\equiv\mathbf{0}$ and another at momentum $\mathbf{k}_{2}\equiv(0,0,\pi)$.
@@
[…]
@@
This Hamiltonian has a single Weyl node remaining at $\textbf{k}_{1}=0$, and no other gapless modes.
@@
Moreover, we see that the BdG generator obeys $S_{\text{chiral}}(\mathbf{k})^{2}=\cos^{2}(k_{z}/2)\mathds{1}$. Thus, it has a continuous spectrum, and is therefore a non-compact symmetry, generating an $\mathbb{R}$ action on the full Hilbert space. This is necessary to evade the no-go theorem in [19], which proved that a locality-preserving chiral $U(1)$ symmetry of a single Weyl fermion does not exist. Curiously, as a Hamiltonian itself, it describes decoupled wires along the $z$-axis of massless Majorana fermions. [^1] A non-Hermitian symmetry generator of a single Weyl fermion was proposed in [20], by interpolating between the identity and a translation symmetry.
We can give an alternative proof of this no-go result, as follows, which applies more generally to the types of symmetries we have been considering. The idea is to observe that $\hat{Q}_{\text{chiral}}$ itself may be viewed as a Hamiltonian with chiral symmetry. If $\hat{Q}_{\text{chiral}}$ could be chosen to have a quantized spectrum, then since it is a two band model, after suitable normalization, one band would be at chiral charge $+1$ and the other at $-1$. As a Hamiltonian then, $\hat{Q}_{\text{chiral}}$ would describe a half-filled band insulator with anomalous chiral $U(1)$ symmetry, which is impossible.
III Symmetry-protected double Weyl fermion in 3+1d
@@
$\begin{gathered}\hat{S}_{0}(\textbf{k})=\tau^{z}\qquad\hat{Q}_{0}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}\hat{S}_{0}(\mathbf{k})d_{\mathbf{k}}\\ \hat{S}_{1}(\textbf{k})=\cos k_{z}\tau^{z}+\sin k_{z}\tau^{x}\qquad\hat{Q}_{1}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}\hat{S}_{1}(\mathbf{k})d_{\mathbf{k}}.\end{gathered}$ (10)
@@
where the identity in the $\tau$ basis ensures it has both $U(1)$ symmetries, $K$ is a parameter, and $m(\textbf{k})$ is the same as in (3). This model is a magnetic Weyl semimetal model that has the two Weyl nodes at $\textbf{k}=\pm\mathbf{K}$ where $\mathbf{K}=(0,0,K)$.
Let’s linearize around the Weyl nodes. We get
@@
$\begin{gathered}\hat{S}_{1,K}(\textbf{k})=\cos K\ \tau^{z}+\sin K\ \tau^{x}.\end{gathered}$ (13)
@@
[…]
@@
IV Time Reversal Symmetric Single Weyl Fermion in 3+1d

### v3

@@ I Introduction
Regulating chiral gauge theories like the Standard Model on the lattice has been a long standing problem. As well as offering a route for extracting more precise predictions from these theories, it also highlights a deep theoretical problem thanks to early no-go results like the Nielsen-Ninomiya theorem [1, 2, 3].
This theorem in its basic form says that in any lattice system of fermions with an on-site $U(1)$ symmetry, that symmetry must act non-chirally in the IR. In particular, if the model has only charge 1 Weyl fermions in the IR, then they must come in equal numbers of left-handed and right-handed chiralities. This is known as fermion doubling.
For quadratic/non-interacting/free Hamiltonian models, we can give an intuitive reason for this phenomenon: Weyl fermions appear in the Brillouin zone as monopoles for the Berry connection [4, 5]. Their total number must therefore be zero because the Brillouin zone is compact and the Berry magnetic field lines form closed loops, just like ordinary magnetic fields.
Fermion doubling remains a problem even if we want to construct chiral gauge theories with vanishing gauge anomalies, such as the Standard Model, so long as we insist on using free fermions. See Appendix A.
There have been several approaches so far to circumvent the fermion doubling problem. A simple strategy is to try to ``gap out'' the unwanted fermions by introducing carefully chosen mass terms. Examples of this construction are Wilson fermions [6] and Kogut-Susskind fermions [7].
As a consequence of the Nielsen-Ninomiya theorem, these mass terms necessarily break any on-site symmetries which would otherwise become chiral symmetries. This puts the fermions in danger of being fine tuned, unless there is enough remnant symmetry to protect them from obtaining other mass terms (i.e. to prevent additive mass renormalization). For example there could be discrete crystalline symmetries of the lattice model such as translation, which can act chirally. This mechanism stabilizes gapless fermions in condensed matter systems such as Weyl semimetals [8, 9, 10, 11, 12].
Another method to stabilize such fermions is to use more general not-on-site symmetries, which can give rise to continuous chiral symmetries. An early example is the Ginsparg-Wilson symmetry of a Euclidean lattice fermion [13, 14] which as an infinitesimal transformation takes the form
@@
where $a$ is the lattice spacing, and $D$ is a finite difference operator which is a discretization of the spacetime Dirac operator. The operator $(D\psi)(x)$ is a (linear) function of $\psi$ at nearby spacetime lattice points, which makes it not-on-site, whereas an on-site symmetry would be a function of $\psi(x)$ alone, such as the vector symmetry. When $a\to 0$ this becomes the continuum axial symmetry of a Dirac fermion. It thus protects the Dirac fermion from gaining a mass. To emphasize that this axial symmetry is not just emergent, but that it is forced by the UV Ginsparg-Wilson symmetry, we call it an ``emanant symmetry'' after [15].
In this paper, we will construct Hamiltonian models with minimal number of fermions at low energy, transforming anomalously under not-on-site symmetries. In particular, we construct two main models:
@@
    An ultralocal [^1] Ultralocal means finite range in real space. 3+1D model with a single Weyl fermion in the IR, protected by an emanant $U(1)$ chiral symmetry which is also ultralocally generated but not quantized in the UV
@@
    An ultralocal 3+1D model with a pair of Weyl fermions, protected by an emanant $SU(2)$ chiral symmetry with Witten's global $SU(2)$ anomaly [16]. The $SU(2)$ has a pair of ultralocal generators which are quantized but which do not close to form $su(2)$, instead generating an infinite-dimensional Lie algebra known as the Onsager algebra.
We emphasize that these models have the minimal number of fermions for their emanant anomalous symmetries, and that each is protected from gaining a mass—they are not fine tuned. In the Appendix, we consider models with time reversal symmetry in 2+1D and 3+1D, but have to relax ultralocality.
Other Hamiltonian families of models of fermions enjoying Ginsparg-Wilson symmetries exist, such as overlap fermions [17, 18, 19, 20, 21]. However, these models are not ultralocal, and can be difficult to compute with. As far as we know, our constructions 1 and 2 give the first ultralocal models with these minimal representations and emanant continuous symmetry groups.
Furthermore, our methods are rather simple, and open a new approach to constructing ultralocal Hamiltonian lattice models with Ginsparg-Wilson-like symmetries. From the condensed matter point of view, one can interpet our models as Weyl semimetals [8, 9]. Using band theory, and especially the Bogoliubov-de-Gennes (BdG) formalism, we are able to study a large class of transformations generalizing the Ginsparg-Wilson symmetry (1) which are linear in the fermion operators, which we leverage to produce the models. We can also derive some no-go results in this formalism, constraining what these not-on-site symmetries must look like for certain anomalous symmetries to emanate from them, making contact with other no-go theorems generalizing Nielsen-Ninomiya such as in [22].
@@ II Symmetry-protected single Weyl fermion in 3+1D
Here we build a (time-reversal broken) 3+1D tight-binding model with finite-range hopping and a single Weyl node at crystalline momentum $\mathbf{k}=\mathbf{0}$ that is protected by a finite-range non-on-site chiral symmetry. The Hamiltonian and symmetry we will construct both lack the full cubic symmetry of the lattice and only retain the $\frac{\pi}{2}$ rotational symmetry around the $\hat{z}$ axis. In Appendix B we present a single symmetry-protected Weyl fermion with the full cubic symmetry. However, we first present the model below because it is much simpler and will inspire the construction of Section IV.
We start with a two-band model (two fermion species per site, which we think of as spin $s\in\{\uparrow,\downarrow\}$) on a cubic lattice, known as a magnetic Weyl semimetal, described by the second-quantized Hamiltonian
@@
with $m(\mathbf{k})={{\Delta}}(2-\cos k_{x}-\cos k_{y})$ for $\Delta\in\mathbb{R}$, and $\boldsymbol{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are Pauli matrices acting on the spin degree of freedom. It hosts two Weyl fermions at low energy when $|\Delta|>1$, seen by linearizing the Hamiltonian are the node at momentum $\mathbf{k}_{1}\equiv\mathbf{0}$ and another at momentum $\mathbf{k}_{2}\equiv(0,0,\pi)$.
@@
The Pauli matrices $\tau^{x,y,z}$ act on a fictitious doubling degree of freedom which separately labels particles at k and holes at $\textbf{k}$ (this is also where the factor of $\frac{1}{2}$ comes from). Thus, any valid Hamiltonian or symmetry generator $h^{\mathrm{BdG}}(\mathbf{k})$ in this BdG formalism must satisfy a particle-hole redundancy
@@
This Hamiltonian has a single Weyl node remaining at $\textbf{k}_{1}=0$, and no other gapless modes, when $|\Delta|>2$.
@@
Moreover, we see that the BdG generator obeys $S_{\text{chiral}}(\mathbf{k})^{2}=\cos^{2}(k_{z}/2)\mathds{1}$. Curiously, as a Hamiltonian itself, it describes decoupled wires along the $z$-axis of massless Majorana fermions. [^2] A non-Hermitian symmetry generator of a single Weyl fermion was proposed in [23], by interpolating between the identity and a translation symmetry. Thus, it has a continuous spectrum, and is therefore a non-compact symmetry, generating an $\mathbb{R}$ action on the full Hilbert space.
In fact, a non-quantized spectrum is necessary to evade the no-go theorem of Fidkowski and Xu [22], a generalization of Nielsen-Ninomiya. Their theorem states that a chiral $U(1)$ symmetry with an exponentially-local charge operator cannot act non-trivially on single Weyl fermion in the IR [^3] This was recently emphasized and explored in [21].. We state and prove a closely related no-go theorem below.
III A No-Go Theorem for $U(1)$ Symmetries
We give a no-go theorem along the lines of Fidkowski and Xu [22] in the context of free-fermion symmetries in arbitrary dimensions.
Let $\hat{Q}_{\text{chiral}}$ be a generator defining a $U(1)$ symmetry (i.e. with quantized charges), and assume that $\hat{Q}_{\text{chiral}}$ is a free-fermion symmetry, and is generated by a translation-invariant single-particle Hamiltonian $S_{\text{chiral}}$ with exponential decay in real space. We will show $\hat{Q}_{\text{chiral}}$ admits a symmetric, gapped, exponentially-local Hamiltonian which is non-degenerate on a torus.
First we observe that $\hat{Q}_{\text{chiral}}$ itself may be viewed as a Hamiltonian with chiral symmetry. Indeed, it is Hermitian and satisfies $\hat{Q}_{\text{chiral}},\hat{Q}_{\text{chiral}}]=0$. By our assumption of the $U(1)$ symmetry, $\hat{Q}_{\text{chiral}}$ has an integer spectrum, so $S_{\text{chiral}}$ does as well. If $S_{\text{chiral}}$ has no zero eigenvalue, then the ground state of $\hat{Q}_{\text{chiral}}$ is the state with all negative eigenvalues of $S_{\text{chiral}}$ filled and all positive eigenvalues of $S_{\text{chiral}}$ empty, with an energy gap to the first positive eigenvalue. This is a band insulator and is non-degenerate on a torus and in this case is what we wanted to show.
Suppose on the other hand that $S_{\text{chiral}}$ has a zero eigenvalue. Then $\hat{Q}_{\text{chiral}}$ has an extensively degenerate flat band in its ground state, with a gap to its further excited states, and we cannot immediately conclude as we did above. Consider the single-particle operator $P(\textbf{k})$ which in the eigenbasis of $S_{\text{chiral}}(\textbf{k})$ acts as $1$ on the zero eigenvectors of $S_{\text{chiral}}(\textbf{k})$ and as 0 on the other eigenvectors. The single-particle Hamiltonian $h(\textbf{k})=S_{\text{chiral}}(\textbf{k})+P(\textbf{k})$ has no zero energy band and commutes with $S_{\text{chiral}}(\textbf{k})$. In real space, it is at least exponentially localized since $S_{\text{chiral}}$ has a uniform band gap of 1 (see Appendix E). Therefore, it defines a local Hamiltonian $\hat{H}$ with $\hat{Q}_{\text{chiral}}$ symmetry. Since we have removed the zero energy band, the ground state of $\hat{H}$ is a gapped band insulator as above, finishing the argument.
By standard anomaly-matching lore, having such a ``trivial'' symmetric Hamiltonian means that $\hat{Q}_{\text{chiral}}$ is free of 't Hooft anomalies. It therefore cannot realize the chiral symmetry of a single Weyl fermion, by anomaly matching.
IV Symmetry-protected double Weyl fermion in 3+1D
@@
$\begin{gathered}S_{0}(\textbf{k})=\tau^{z}\qquad\hat{Q}_{0}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}S_{0}(\mathbf{k})d_{\mathbf{k}}\\ S_{1}(\textbf{k})=\cos k_{z}\tau^{z}+\sin k_{z}\tau^{x}\qquad\hat{Q}_{1}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}S_{1}(\mathbf{k})d_{\mathbf{k}}.\end{gathered}$ (10)
@@
where the identity in the $\tau$ basis ensures it has both $U(1)$ symmetries, $K$ is a parameter, and $m(\textbf{k})$ is the same as in (3). When $|\cos K-\Delta|>1$, this model is a magnetic Weyl semimetal model with two Weyl nodes of opposite chirality at $\textbf{k}=\pm\mathbf{K}$ where $\mathbf{K}=(0,0,K)$.
Let's linearize around the Weyl nodes. We get
@@
$\begin{gathered}S_{1,K}(\textbf{k})=\cos K\ \tau^{z}+\sin K\ \tau^{x}.\end{gathered}$ (13)
@@
Thus, together with $S_{0}=\tau^{z}$, these generate an $su(2)$ algebra acting on the low energy theory. Note that $\tau^{x}$ acts by exchanging particles at K with holes at $\textbf{K}$. Thus, it is convenient to apply a charge conjugation the right-handed Weyl fermion, to give a low energy theory in terms of two left-handed Weyl fermions, now with opposite charge w.r.t. $\tau^{z}$ rotations. $\tau^{x}$ rotations meanwhile act by a flavor rotation exchanging the two Weyl fermions. Thus, our symmetry generators $\hat{Q}_{0}$, $\hat{Q}_{1}$ correspond to two $su(2)$ generators in the flavor symmetry of the low energy, at an angle of $K$. For $K\neq 0,\pi$, they thus generate the whole chiral symmetry.
@@
V Discussion
The chiral symmetry in 3+1D has a nearest-neighbor charge operator, but is non-quantized, as it must be by the no-go theorem of [22] and our arguments above. The two $SU(2)$ generators we constructed are quantized (and one is on-site), but do not satisfy the expected $SU(2)$ Lie algebra, instead generating an infinite-dimensional Onsager algebra. This is consistent with the anomaly since either $U(1)$ symmetry on its own is anomaly-free. In this way, it is structurally similar to the $U(1)_{V}\times U(1)_{A}$ anomaly realized in 1+1D in [27, 28], where both generators are quantized and one is on-site, but they don't commute.
There is a fun example which avoids this no-go argument in 2+1D. Let us take a 2d square lattice of spin-$\frac{1}{2}$ degrees of freedom, and take $\hat{Q}$ to be the toric code Hamiltonian [29], which has an integer spectrum. It also commutes with time reversal given by complex conjugation, and together, these two symmetries generate a bosonic $U(1)\rtimes T$ parity anomaly [30]. The ground state of the toric code Hamiltonian is indeed non-trivial (it is degenerate on a torus), and sufficient to match this parity anomaly.
Finally, we would eventually like to make lattice models of chiral gauge theories. Although we have produced models with chiral symmetries, the not-on-siteness of the symmetries makes them difficult to gauge. In fact one expects on general grounds that because of the nonvanishing 't Hooft anomalies, we should not be able to gauge these symmetries. For interesting anomaly cancellation, such as in the Standard Model, one needs low energy fermions of different charges. It is not clear models of the kind we studied can produce such charge assigments. It would be very interesting if in some more complicated models with not-on-site chiral symmetries having vanishing 't Hooft anomalies, if we can nonetheless find some way to gauge them, as we done for discrete symmetries in 1+1D in [31].
Appendix A Fermion Doubling and Anomalies
We can also reason about fermion doubling from the point of view of 't Hooft anomaly matching [32]. In particular, a system with an on-site global symmetry, meaning one which does not mix degrees of freedom at separate spacetime points, is free of 't Hooft anomalies in the UV, since a 't Hooft anomaly is an obstruction to gauging and there is a canonical method to gauge on-site symmetries. By anomaly matching, whatever theory arises from this lattice model in the IR must also be free of 't Hooft anomalies.
For an IR composed of $N_{L}$ left-handed Weyl fermions of charges $q_{i}^{L}$ and $N_{R}$ right-handed Weyl fermions of charge $q_{i}^{R}$, the anomaly-vanishing equations are
$\begin{gathered}\sum_{i=1}^{N_{L}}q_{i}^{L}=\sum_{i=1}^{N_{R}}q_{i}^{R}\qquad\sum_{i=1}^{N_{L}}(q_{i}^{L})^{3}=\sum_{i=1}^{N_{R}}(q_{i}^{R})^{3}.\end{gathered}$ (15)
Trivial solutions are those with $N_{L}=N_{R}$ and $q_{i}^{L}=q_{i}^{R}$ where fermions come in non-chiral pairs. For chiral gauge theories like the Standard Model, we want the IR to realize a non-trivial solution to these equations.
However, in a free-fermion (quadratic) Hamiltonian model, fermions of different charge never interact, and so the anomaly-vanishing equation has to hold for each $U(1)$ charge separately, yielding only trivial solutions with fermion doubling.
Appendix B Cubic-symmetric single Weyl fermion model
In this appendix we will present a model of a symmetry-protected single Weyl fermion that is fully cubic-symmetric, as opposed to the the model in Section II. Here, we will take the BdG basis $d_{\mathbf{k}}^{\dagger}=(c_{\mathbf{k}}^{\dagger},[i\sigma^{y}c_{-\mathbf{k}}]^{T})$, which imposes the particle-hole redundancy
$\tau^{y}\sigma^{y}h^{\mathrm{BdG}}(\mathbf{k})^{T}\tau^{y}\sigma^{y}=-h^{\mathrm{BdG}}(-\mathbf{k})\quad,$ (16)
as opposed to Eq. 5.
Consider the Bloch Hamiltonian in the BdG formalism given by
$h^{\rm BdG}(\mathbf{k})=\sin k_{x}$ $\tau^{z}\sigma^{x}+\sin k_{y}\tau^{z}\sigma^{y}$
$+\sin k_{z}\tau^{z}\sigma^{z}+M(\mathbf{k})\tau^{x}\quad,$ (17)
where $M(\mathbf{k})=3-\cos k_{x}-\cos k_{y}-\cos k_{z}$. This model possesses a single massless Weyl fermion at $\mathbf{k}=0$, and retains the cubic symmetry of the lattice, with the $\frac{\pi}{2}$ rotations given by $i\sigma^{x,y,z}$ as well as time-reversal symmetric. A similar model was given in the original Nielsen-Ninomiya paper [3].
We can construct a chiral symmetry generator $S_{\rm chiral}(\mathbf{k})$ that protects this massless Weyl fermion, with $S_{\rm chiral}(\mathbf{k})$ given by
$S_{\rm chiral}(\mathbf{k})=$ $\frac{1}{8}\big[(1+\cos k_{x})(1+\cos k_{y})(1+\cos k_{z})\tau^{z}$
$+\sin k_{x}(1+\cos k_{y})(1+\cos k_{z})\tau^{x}\sigma^{x}$
$+(1+\cos k_{x})\sin k_{y}(1+\cos k_{z})\tau^{x}\sigma^{y}$ (18)
$+(1+\cos k_{x})(1+\cos k_{y})\sin k_{z}\tau^{x}\sigma^{z}\big],$
where $S_{\rm chiral}(\mathbf{k}),h^{\rm BdG}(\mathbf{k})]=0$. Observe that
$S_{\rm chiral}(\mathbf{k})=\begin{cases}\tau^{z}&{\rm when}\,\,\mathbf{k}=\mathbf{0}\\ 0&{\rm when}\,\,k_{x}=\pi\\ 0&{\rm when}\,\,k_{y}=\pi\\ 0&{\rm when}\,\,k_{z}=\pi\end{cases}\quad,$ (19)
i.e., $S_{\rm chiral}(\mathbf{k})$ acts like the usual $U(1)$ symmetry $\tau^{z}$ at $\mathbf{k}=0$, where the massless Weyl fermion is, but is 0 at the boundary of the Brillouin zone. This symmetry is still unquantized, i.e. the charge generator is not quantized. The physical interpretation of such a symmetry, when Fourier transformed back to real space, is not as simple the bundle of critical Kitaev chains in the example given in Eq. 7 of the main text. In this case, the real space Hamiltonian would involve fermion hoppings of range three, whose ground state corresponds to a Bogoliubov Fermi surface state.
The Hamiltonian in Eq. 17 with the chiral symmetry, generated by $S_{\rm chiral}(\mathbf{k})$, still possesses terms that can eventually (non-perturbatively) gap the single massless Weyl fermion since the system does not possess a quantized UV anomaly  [^4] A more detailed handling of quantized versus unquantized anomalies can be found in Ref. [33].. $S_{\rm chiral}(\mathbf{k})$ is itself such a term – upon adding it with a magnitude $\delta>0$ to Eq. 17, the Weyl fermion at the origin transitions into a Fermi surface with the low-energy linearization around $\mathbf{k}=\mathbf{0}$ being
$h^{\rm BdG}(\mathbf{k})\approx k_{x}$ $\tau^{z}\sigma^{x}+k_{y}\tau^{z}\sigma^{y}+k_{z}\tau^{z}\sigma^{z}+\delta\tau^{z}\quad,$ (20)
when $\delta$ is small. When the magnitude $\delta$ is increased, this Fermi surface expands and then eventually gaps at the boundary of the Brillouin zone.
Appendix C Time Reversal Symmetric Single Weyl Fermion in 3+1D
So far, we have considered time-reversal breaking models which are ultralocal in space, having bounded-range hoppings, with symmetries satisfying the same. We can also construct time-reversal invariant models, but this will come at the cost of making the models and their symmetries non-ultralocal. We do not have a proof that it is necessary to do this, and in the future it would be interesting to explore whether it is necessary is not.